|
Документ взят из кэша поисковой машины. Адрес
оригинального документа
: http://www.stsci.edu/smov/notes/0303smov.html
Дата изменения: Tue Mar 11 23:59:19 1997 Дата индексирования: Thu Feb 27 23:54:31 2014 Кодировка: Поисковые слова: http www.astronomy.com |
STATUS REPORTS
1. Observatory Status
MOSES status reports:
-PCS - Nominal
- Gyro biases continue to be monitored.
- The first STIS acquisition was monitored closely
in case of any acquisition problems. There were
none.
- RWA1 is being monitored during slews, especially
now that we are out of the BEA period.
- Other observatory subsystems all nominal
2. WFPC2 Status
- Engineering status is nominal. A DECON is scheduled
for tonight.
- Science status - nominal. UV throughput continues to
be monitored with no significant difference from
pre-SM2 measurements.
3. NICMOS Status
- Engineering status - NICMOS experienced another
transition to SUSPEND state this morning at 12:32 UT.
The event appears to have the same signature as the
previous transition to SUSPEND, occurring again deep
in the SAA. This time, however, it occurred on
Detector 3 instead of 1. There will be a meeting later
this morning to firm up plans for any realtime manipula-
tions that need to preceed the scheduled Pre-Alignment
Check-out (7134) scheduled to occur tomorrow.
There is also a 5:30 pm (EST) meeting
today to discuss the NICMOS SAA Contours Test, scheduled
farther downstream in SMOV.
- Science Status - nominal - Data analysis continues on
previously executed internal activities. A report on
part 2 of the Parallel Operations Test (7136) will be
out soon.
4. STIS Status
- Engineering Status - Nominal - The MAMA LVPSs remain
off indefinitely, pending further investigation by BASD.
The Ball ground testing has produced a few instances of HV
turn-on while simulating the in-orbit conditions for the
MCE reset anomalies. Life testing of the Band2 flight spare
is being conducted in order to determine safe ramp
times as a potential operational implementation for MAMA
high-voltage operations. The STIS bulkhead is running
several degrees cooler than anticipated because the MAMA
LVPSs are off. Meetings tomorrow and, if necessary, through
the rest of the week, will determine if the SMOV schedule
is impacted in any way by these problems.
- Science Status - STIS reports that the first STIS
external observation, a star field viewed by the CCD
for the purpose of STIS/FGS alignment (7065) is a success.
The target appears well centered in the FOV with a FWHM
of less than 2 pixels and only a slight hint of coma.
Data analysis is still in the early stages but the STIS
astronomers who have seen the image are quite pleased.
- The STIS team also report a red leak in the MgII filter
and a pinhole in the 270 continuum filter. In
addition the fiducial (0.2 X 0.2) aperture displays a small
amount of orbital motion (0.15 pixels), attributable to
STIS heater cycling. These problems are all characterized
as minor.
5. FGS status - no activity
The first FGS1R activity (7026, visit 1) is scheduled
for later this morning.
6. FOC/COSTAR status - nominal - no activity
The FOC non-COSTAR baseline throughput measurements (7012)
are scheduled for this evening.
PLANNING REPORTS
1. Status of Upcoming SMSs
- SMS062 - Currently executing.
- SMS065 - Signed off, rady for uplink.
- SMS069 - Finals are out for review.
- SMS072 - Built and initials out for review
- SMS076 - Built and initials out for review
- SMS079 - Initial build complete. This SMS currently
contains the 1st part of the MAMA HV turn-on,
per the nominal plan.
2. Replans - none
OPERATIONS STATUS
1. OPUS - A fraction of STIS data is not making it through
calibration, partly due to an IRAF kernal problem that
will be corrected this week, and partly due to misconfigured
flags and reference files in the pipeline. All problems
have operational workarounds in place and permanent
corrections planned.
2. Archive - Nominal